# UStyle: Waterbody Style Transfer of Underwater Scenes by Depth-Guided Feature Synthesis

### v2

@@ II-D Underwater Domain Style Transfer
Traditionally, domain STx methods are used to model underwater image distortions for synthetic data generation [93, 94]; the stylized distorted pairs are then used for learning image enhancement [95, 11] or improved object detection [15, 96]. These models enhance color correction, clarity, and stylization in a unified learning pipeline by separating content and style across different domains. For instance, Chen et al. [47] introduces a domain adaptation framework that leverages transformers for image enhancement [97, 98]. Zhou et al. [99] uses a transformer-based STx model for stylized representation learning of target recognition in underwater acoustic imagery. Our previous work AquaFuse [1] proposes a physics-guided method for geometry-preserving waterbody fusion for photometric data augmentation, which requires a prior reference image and known waterbody parameters. This work explores no-reference image STx from comprehensive multimodal learning databases – for the first time in the literature.
Figure 3: A few samples from the UF7D dataset are shown for each of its seven style categories.
@@ III-A Data Preparation and Categorization
To enable data-driven learning of no-reference waterbody STx, we compile a comprehensive collection of high-resolution underwater image database named UF7D. It includes $7$ style categories: Blue (B), Blue-Green (B-G), Clear (C), Deep Blue (DB), Deep Green (DG), Green (G), and Green-Blue (G-B). A few samples for each category are shown in Figure 3.
(a) Distribution of images for all categories.
  Resolution # of Images
  $640\times 480$ $1633$
Up to $800\times 600$ $56$
Up to $1024\times 768$ $106$
Up to $1280\times 720$ $111$
Up to $1920\times 1080$ $374$
Above $1920\times 1080$ $1770$
Total # of Images 4050
(a) Image resolution information.
Figure 4: An overview of the UF7D dataset: it has a balanced distribution of seven waterbody style categories, each consisting of high-resolution natural underwater images.
A key feature of UF7D dataset is that it includes the scene depth maps for each image, generated by using the DepthAnything V2 model [100]. With high-resolution images and depth information, the UF7D dataset supports deep visual learning applications such as underwater image style transfer, enhancement, object detection, and more. To this end, our data collection process focused on the following aspects:
  1. $1$)
    Scene Diversity: UF7D includes a wide variety of natural underwater scenes that are sourced from various benchmark image databases [11, 12, 13, 101, 14, 15] and other public platforms such as Unsplash [102] and Flickr [103]. Our collection spans various objects including coral reefs, shipwrecks, pipelines, and various marine species (fish, plants, rocks) – enabling robust data-driven learning across diverse underwater conditions [104, 105].
  2. $2$)
    Image Quality: UF7D has a total of $4,050$ images, over $44\%$ of which are above $1920\times 1080$ resolution (see Fig. 4), offering high-resolution image contents necessary for fine-grained analyses.
  3. $3$)
    Balanced STx Categories: We relied on human visual perception for curating and compiling the seven style categories. As shown in Figure 4(a), these categories are evenly represented across the dataset, comprising the spectrum of dark blue to dark green colors, as well as clear images.
@@ III-B Dataset Statistics and Analyses
Our statistical analysis reveals that UF7D consists of comprehensive coverage of varying underwater depths. Figure 5(a) illustrates the depth distributions by violin plots, where width represents the frequency density at each depth level. Here, we make the following observations.
@@ 1. $\bullet$
    Cross-category Variation: Median depths range from $60$ to $116$ pixels across categories, with Green (G) environments showing the highest median and Deep Blue (DB) the lowest.
@@ 2. $\bullet$
    Within-category Distribution: Categories exhibit distinct depth patterns; for instance, Clear (C) waters show consistent depth profiles, while Deep Blue (DB) shows a much wider variance.
(a) Violin plots showing depth value distributions across all categories. A particular example of depth variation is shown on the right; the blue box highlights a background region and the red box indicates a foreground region.
(b) PCA projection of global color characteristics (with $92\%$ variance captured in first two components).
(c) t-SNE visualization of local color statistics; the clusters validate the human-perceived category distinctions.
Figure 5: Statistical analyses and relevant information on the UF7D dataset.
We further perform detailed statistical analysis to validate our human-perceived style categorization process. Figure 5 presents the following two complementary analyses:
  1. $\bullet$
    Global Color Structure: Principle component analysis (PCA) reveals distinct clustering patterns for all categories, with the first two principal components capturing $92\%$ of total color variance; see Figure 5(b). The primary component embeds $76\%$ variance, effectively separating blue-dominated waters from green-dominated waters, which is consistent with our human perceptual evaluations.
  2. $\bullet$
    Local Color Relationships: A t-SNE (t-distributed stochastic neighbor embedding) visualization demonstrates clear separations among the style categories; see Figure 5(c). This confirms that images within each category share consistent spatial color characteristics.
These analyses validate that our waterbody style categorization based on human perception is statistically consistent as well. The inter-category and intra-category statistics also suggest that the underlying style can potentially be learned with tailored training strategies.
@@ IV UStyle Model & Learning Pipeline
We present UStyle, a novel waterbody style transfer framework that integrates a deep encoder-decoder architecture, progressive blockwise training, and a depth-aware stylization module. Given a content image $\mathbf{I}_{c}\in\mathbb{R}^{H\times W\times 3}$ and a style image $\mathbf{I}_{s}\in\mathbb{R}^{H\times W\times 3}$, UStyle generates a stylized output $\mathbf{I}_{cs}$ by fusing waterbody characteristics from the style image while preserving scene geometry of the content image. The depth maps $\mathbf{D}_{s}$ and $\mathbf{D}_{c}$, representing the depth information of the style and content images, respectively, adapt features based on depth. $\mathbf{D}_{s}$ guides the extraction of waterbody details from the style image, while $\mathbf{D}_{c}$ ensures content-aware stylization by preserving structural consistency and depth-dependent variations during style transfer. As shown in Fig. 6, the UStyle learning pipeline comprises three major components: (1) an encoder-decoder backbone for multi-scale feature extraction with fine spatial details and high-level semantics; (2) a progressive training strategy for stable coarse-to-fine stylization; and (3) a depth-aware stylization module for effective waterbody style transfer.
Figure 6: Overview of our UStyle model, featuring an encoder-decoder backbone with skip connections, high-frequency residual (HFR) paths, and a CLIP-based coarse similarity module for aligning intermediate features with semantic cues. A multi-stage learning strategy is adopted that progressively trains deeper symmetric blocks to preserve both coarse structure and fine details in the waterbody fusion process.
@@ IV-A Encoder–Decoder Backbone
We use a ResNet50 [18] encoder ($\mathcal{E}$) partitioned into five sequential blocks $\mathrm{EB}_{1:5}$, which extract increasingly finer features $\{{F}_{i,c}\}_{i=0:4}$ and $\{{F}_{i,s}\}_{i=0:4}$ from a given content and style images, respectively. These feature maps are then fed into a decoder $\mathcal{D}$ with hierarchical ($\mathrm{EB}_{i-1}\to\mathrm{DB}_{i}$) skip connections. The symmetric skip connections retain spatial structures and propagate high-frequency details from the encoder to the decoder. Each decoder block $\mathrm{DB}_{i}$ consists of a transposed convolution [106] for upsampling, followed by a convolutional layer [107], batch normalization [108], and a ReLU activation [109]. Rather than explicitly computing residuals via average pooling and subtraction, our design preserves these details through the original skip connections. These features are fused to achieve waterbody STx with pixel-wise guidance from the content depth map $\mathbf{D}_{c}$, ensuring that the fused features ${F}_{i,cs}$ adapt to both structural (object geometry) and stylistic (waterbody) elements. This fusion is performed by our novel depth-aware stylization module; see Section IV-C.
IV-B Progressive Pre-Training and Domain-specific Finetuning
We implement a sequential blockwise approach in UStyle, where the encoder-decoder network is optimized progressively over five stages ($N=1$:$5$). Completed $\mathrm{EB}_{i}-\mathrm{DB}_{i}$ blocks are frozen to ensure stability of the training, as shown in Fig. 6. Initially, the model is pre-trained on a large-scale atmospheric dataset (MSCOCO-2017 [110]) with a weighted combination of reconstruction loss [111], feature reconstruction loss [86], and perceptual loss. During the training of the first stage ($N=0$), the network is optimized using only a reconstruction loss $\mathcal{L}_{r}=\|\mathbf{I}_{c}-\hat{\mathbf{I}}_{c}\|_{2}^{2}$ where $\hat{\mathbf{I}}_{c}$ denotes the decoder’s output. In subsequent stages ($N>0$), additional losses are introduced to ensure feature-level consistency and perceptual fidelity. These include a feature reconstruction loss $\mathcal{L}_{feat}=\|\mathbb{F}_{i}-\hat{\mathbb{F}}_{i}\|_{2}^{2}$ and a perceptual loss $\mathcal{L}_{percept}=\|\phi(\mathbf{I}_{c})-\phi(\hat{\mathbf{I}}_{c})\|_{2}^{2}$, where $\phi(\cdot)$ is a high-level feature extractor based on VGG-16 [24].
After pretraining, UStyle is fine-tuned on a large-scale underwater dataset (USOD10K [112]) for domain-specific supervision with SSIM loss $\mathcal{L}_{ssim}=1-SSIM(\mathbf{I}_{c},\hat{\mathbf{I}}_{c})$, color consistency loss in LAB image space $\mathcal{L}_{color}=\|\mathrm{LAB}(\mathbf{I}_{c})-\mathrm{LAB}(\hat{\mathbf{I}}_{c})\|_{2}^{2}$, and frequency-domain loss by Fast Fourier Transforms (FFT) $\mathcal{L}_{fft}=\|\mathcal{F}(\mathbf{I}_{c})-\mathcal{F}(\hat{\mathbf{I}}_{c})\|_{2}^{2}$. Moreover, CLIP-based content supervision is incorporated by introducing a CLIP loss [25] to ensure semantic alignment. It is defined as:
@@ IV-B Progressive Pre-Training and Domain-specific Finetuning
The overall objective functions for multi-stage training are as follows:
$\mathcal{L}_{0}$ $=\mathcal{L}_{r}+\mathcal{L}_{ssim}+\mathcal{L}_{color}+\mathcal{L}_{fft}+\mathcal{L}_{clip}\quad(N=0)$ (2)
$\mathcal{L}_{N}$ $=\mathcal{L}_{r}+\mathcal{L}_{ssim}+\mathcal{L}_{color}+\mathcal{L}_{fft}+\mathcal{L}_{clip}+\mathcal{L}_{feat}+\mathcal{L}_{percept}\quad(N=1,2,3,4)$ (3)
This progressive training approach allows the network to learn robust multi-scale feature representations while ensuring content, style, and depth consistency for high-fidelity underwater waterbody style transfer.
@@ IV-C Depth-aware Stylization
We design a depth-aware stylization function: DA-WCT which extends the multi-scale blending concept from the WCT (whitening and coloring transform) literature [37]. DA-WCT integrates a physics-guided waterbody estimation method for depth-guided feature blending via WCT. First, we estimate the waterbody from the style image $\mathbf{I}_{s}$ using a physics-guided method [1]; from its depth map $\mathbf{D}_{s}$, we leverage the $5\%$ farthest pixels’ median values to estimate the waterbody background in $\mathbf{I}_{s}$. A color filter is then applied to retain only bluish-green pixels, removing red-channel influence due to differential light attenuation underwater. The extracted waterbody $\mathbf{B}_{s}$ is then fed into the encoder to obtain the style features.
Let $F_{c}$ and $F_{s}$ denote the content and style feature maps extracted from the encoder, respectively. We first perform a WCT-based stylization to align the statistics of $F_{c}$ with those of $F_{s}$; the stylized feature $\tilde{F}_{cs}$ is computed as $\tilde{F}_{cs}\;=\;\mu_{s}\;+\;C_{s}^{1/2}\,C_{c}^{-1/2}\,\small(F_{c}-\mu_{c}\small)$, with $\mu_{c}$, $\mu_{s}$ being the channel-wise means and $C_{c}$, $C_{s}$ the covariance matrices of $F_{c}$ and $F_{s}$, respectively. Let $\mathbf{D}_{c}$ denote the single-channel content depth map, a global depth threshold $\tau\in[0,1]$ is computed using Otsu’s method [113]. Next, an adaptive scaling factor $k$ is derived by first smoothing $\mathbf{D}_{c}$ via average pooling with a $5\times 5$ kernel, then computing the global mean and standard deviation as follows:
$\mu_{d}$ $=\frac{1}{H\cdot W}\sum_{x}\mathrm{avgpool}\big{(}\mathbf{D}_{c}(x)\big{)}$ (4)
@@
$\sigma_{d}$ $=\sqrt{\frac{1}{H\cdot W}\sum_{x}\Bigl{(}\mathrm{avgpool}\big{(}\mathbf{D}_{c}(x)\big{)}-\mu_{d}\Bigr{)}^{2}}$ (5)
@@
With the tunable scaling factors $\tau$ and $k$ determined, the depth weight map for each pixel is computed using the sigmoid function.
@@
$w(x)=\sigma\Bigl{(}-k\cdot\bigl{(}\mathbf{D}_{c}(x)-\tau\bigr{)}\Bigr{)}=\frac{1}{1+\exp\Bigl{(}k\cdot\bigl{(}\mathbf{D}_{c}(x)-\tau\bigr{)}\Bigr{)}}$ (6)
Subsequently, the blended feature is obtained by merging the stylized features $\tilde{F}_{cs}(x)$ with the original content features $F_{c}(x)$ as follows
$F_{\text{depth}}(x)=w(x)\cdot\tilde{F}_{cs}(x)+[1-w(x)]\cdot F_{c}(x)$ (7)
@@
In our formulation, pixels with higher depth values (i.e., farther objects) yield $w(x)\rightarrow 1$ and favor the stylized features $\tilde{F}_{cs}(x)$. In contrast, pixels with lower depth values get $w(x)\rightarrow 0$, preserving more of the original content features $F_{c}(x)$. A global style strength parameter $\alpha\in[0,1]$ is applied to control the degree of the underlying stylization process.
$F_{cs}\;=\;\alpha\,F_{\text{depth}}\;+\;\bigl{(}1-\alpha\bigr{)}\,F_{c}.$ (8)
[…]
$\mathbf{I}_{stylized}=\text{GIF}(\mathbf{I}_{cs})=\text{GIF}\big{(}\mathcal{D}(F_{\text{multi}})\big{)}.$ (9)
@@
Figure 7: Inference pipeline of the proposed UStyle model. The content and style images and their corresponding depth maps are first processed through a ResNet-based encoder to extract multi-scale features. These features are processed by our depth-aware whitening and coloring transform (DA-WCT) function for waterbody style alignment, which are then progressively decoded using a multi-stage decoder. Finally, the stylized output is refined with a guided filter to improve visual consistency.
@@
Implementation Details. UStyle is implemented in PyTorch with OpenCV and Torch libraries. We use Adam optimizer with an initial learning rate of $10^{-4}$ for the supervised learning. Training is conducted on a Linux system equipped with $4$ GPUs and $32$ GB of RAM. At the first stage of training, the encoder-decoder network is pre-trained on the MSCOCO-2017 with $118$ K training samples for $10$ epochs. It is then fine-tuned on USOD10K underwater dataset with $10$ K examples for another $10$ epochs. The whole training and fine-tuning takes $82$ hours to complete. Once trained, the frozen graph is used for inference, executed on a single GPU and $16$ GB of RAM.
Inference Pipeline. The inference process of UStyle follows a structured pipeline to achieve high-fidelity underwater style transfer while maintaining the structural integrity of the content image. As illustrated in Fig. 7, the process begins with a given content image $\mathbf{I}_{c}$ and a style image $\mathbf{I}_{s}$, along with their corresponding depth maps $\mathbf{D}_{c}$ and $\mathbf{D}_{s}$. The style image is first processed by our physics-guided Waterbody Extractor [1], which isolates the dominant waterbody characteristics $\mathbf{B}_{s}$. Both $\mathbf{I}_{c}$ and $\mathbf{B}_{s}$ are then passed into the Encoder, which extracts multi-scale feature representations $\mathbf{F}_{c}$ and $\mathbf{F}_{s}$, respectively. The extracted features are then processed by our DA-WCT module, which aligns the feature statistics between content and style domains, producing stylized feature representations $\mathbf{F}_{cs}$. These transformed features are fed into a multi-stage Decoder, each progressively refining the stylized output. Skip connections from the Encoder to the Decoder help retain spatial consistency and fine-grained structural details. At each decoding stage, an adaptive fusion parameter $\alpha$ controls the style influence level, ensuring smooth content-style integration. Finally, the stylized image $\mathbf{I}_{cs}$ is passed through a Guided Filter, which enhances structural coherence and removes unwanted artifacts, yielding the final stylized output $\mathbf{I}_{stylized}$.

### v3

@@ II-D Underwater Domain Style Transfer
Traditionally, domain STx methods are used to model underwater image distortions for synthetic data generation [fabbri2018enhancing, li2017watergan]; the stylized distorted pairs are then used for learning image enhancement [li2018synthesis, li2019underwater, zhang2023underwater, zhuang2024globally, wu2024underwater, zhou2025net, sun2022underwater, zhou2023underwater, zhang2022underwater, wang2023underwater, hu2023enhancing] or improved object detection [islam2020fast, yang2022underwater]. These models enhance color correction, clarity, and stylization in a unified learning pipeline by separating content and style across domains. Chen et al. [chen2022domain] introduces a domain adaptation framework that leverages transformers for image enhancement [anil2023novel]. Zhou et al. [zhou2023learning] uses a transformer-based model for stylized representation learning of target recognition in underwater acoustic imagery. Our previous work AquaFuse [siddique2024aquafuse] proposes a physics-guided method for geometry-preserving waterbody fusion for data augmentation, which requires a prior reference image and known waterbody parameters. We address these knowledge gaps by exploring no-reference waterbody STx from comprehensive multimodal databases.
(a) Sample images and corresponding depth maps for each style category.
(b) Distribution of the seven style categories in UF7D.
(c) PCA projection of global color characteristics ($92\%$ variance).
(d) t-SNE visualization of image statistics confirming category distinctions.
Figure 3: Sample images and statistical analyses of the UF7D dataset.
@@ III-A Data Preparation and Categorization
We present UF7D, a high-resolution underwater image dataset comprising seven waterbody styles: Blue (B), Blue-Green (B-G), Clear (C), Deep Blue (DB), Deep Green (DG), Green (G), and Green-Blue (G-B). UF7D combines both benchmark-sourced and newly collected data to ensure broad diversity across underwater environments. Approximately $58\%$ of the images come from existing datasets (UIEB [li2019underwater], Sea-Thru [akkaynak2019sea], LSUI [peng2023u], U45 [li2019underwater], EUVP [islam2020fast]), while the other $42\%$ are collected from community and open-access sources (e.g., Unsplash, Flickr), with the focus to introducing diverse scenes and/or waterbody types that are not common in prior benchmarks. The dataset spans varied geographic regions and water types, covering coral reefs, marine species, shipwrecks, pipelines, rocks, and other natural structures. For each image, a relative depth map is generated using DepthAnythingV2 [yang2024depth].
UF7D dataset contains $4,050$ images, with over $44\%$ exceeding $1920\times 1080$ resolution, enabling fine-grained learning. There are 25 test samples per category, with $625$ content-style evaluation pairs; representative samples and depth maps are shown in Fig. 3(a). To further assess STx generalization, we release a UF7D-Challenge test set with $70$ images across categories, which captures a set of challenging regional, temporal, and sensory variations. The images were collected from online open-access sources and in-field captures, covering diverse locations, lighting, and sensors. This enables testing under unseen environmental and acquisition conditions, following recent evaluation protocols [Li_Hyunida_2024, boosting_li_2025, LIANG2025104463].
@@ III-B Dataset Statistics and Analyses
UF7D combines benchmark-sourced images with a substantial portion of newly collected, high-resolution data, forming a diverse and well-balanced dataset across different water types and scene compositions. To validate the consistency and distinctness of its seven waterbody categories, we perform global and local statistical analyses as follows:
@@ 1. $\bullet$
    Global color structure: PCA shows clear separation among the seven waterbody styles, with the first two components explaining $92\%$ of color variance; see Fig. 3(c). The dominant component ($76\%$ variance) separates blue- and green-dominated waters, validating the perceptual categorization.
@@ 2. $\bullet$
    Local statistics: The t-SNE projections shown in Fig. 3(d) reveal distinct, non-overlapping clusters, indicating consistent intra-category characteristics and strong inter-category differences.
These statistics confirm that UF7D is both perceptually and statistically well balanced, with sufficient unseen content to support robust, data-driven learning of underwater styles.
@@ IV UStyle Model & Learning Pipeline
We present UStyle, a novel waterbody style transfer framework that integrates a deep encoder-decoder architecture, progressive blockwise training, and a depth-aware stylization module. Given a content image $\mathbf{I}_{c}\in\mathbb{R}^{H\times W\times 3}$ and a style image $\mathbf{I}_{s}\in\mathbb{R}^{H\times W\times 3}$, UStyle generates a stylized output $\mathbf{I}_{cs}$ by fusing waterbody characteristics from the style image while preserving scene geometry of the content image. The depth maps $\mathbf{D}_{s}$ and $\mathbf{D}_{c}$, representing the depth information of the style and content images, respectively, adapt features based on depth. $\mathbf{D}_{s}$ guides the extraction of waterbody details from the style image, while $\mathbf{D}_{c}$ ensures content-aware stylization by preserving structural consistency and depth-dependent variations during style transfer. The detailed architecture and learning pipeline is outlined in Fig. 4.
@@ IV-A Encoder–Decoder Backbone
We employ a ResNet50 [koonce2021resnet] encoder ($\mathcal{E}$) with five blocks $\mathrm{EB}_{1:5}$ to extract features $\{F_{i,c}\}_{i=0:4}$ and $\{F_{i,s}\}_{i=0:4}$ from the content and style images. These features are passed to a decoder $\mathcal{D}$ via hierarchical skip connections to retain spatial details. Each decoder block $\mathrm{DB}_{i}$ uses a transposed convolution [gao2019pixel] for upsampling, followed by convolution, batch normalization, and ReLU activation. Guided by the content depth map $\mathbf{D}_{c}$, features are fused into ${F}_{i,cs}$—capturing both structural (object geometry) and stylistic (waterbody) elements—via our depth-aware stylization module; see Section IV-C.
Figure 4: Overview of our UStyle model, an encoder-decoder with skip connections, high-frequency residual (HFR) paths, and a CLIP-based module aligning features with cues. Multi-stage training progressively trains deeper symmetric blocks to preserve both coarse structure and fine details in waterbody fusion.
IV-B Pre-Training and Domain-specific Finetuning
We adopt sequential blockwise training in UStyle over five stages ($N$), freezing incremental $\mathrm{EB}_{i}-\mathrm{DB}_{i}$ blocks for stability. Initially, the model is pre-trained on a large-scale atmospheric dataset (MSCOCO-2017 [lin2014microsoft]) with a weighted combination of reconstruction loss [rumelhart1986learning], feature reconstruction loss [johnson2016perceptual], and perceptual loss. At stage NN=0, only the reconstruction loss is used: $\mathcal{L}_{r}=\|\mathbf{I}_{c}-\hat{\mathbf{I}}_{c}\|_{2}^{2}$, where $\hat{\mathbf{I}}_{c}$ is the decoder output. Subsequently, additional feature loss $\mathcal{L}_{feat}=\|\mathbb{F}_{i}-\hat{\mathbb{F}}_{i}\|_{2}^{2}$ and perceptual loss $\mathcal{L}_{percept}=\|\phi(\mathbf{I}_{c})-\phi(\hat{\mathbf{I}}_{c})\|_{2}^{2}$ are added where $\phi(\cdot)$ is a content loss based on VGG-16 [simonyan2014very].
After pretraining, UStyle is fine-tuned on a large-scale underwater dataset (USOD10K [hong2023usod10k]) for domain-specific supervision with SSIM loss $\mathcal{L}_{ssim}=1-SSIM(\mathbf{I}_{c},\hat{\mathbf{I}}_{c})$, color consistency loss in LAB image space $\mathcal{L}_{color}=\|\mathrm{LAB}(\mathbf{I}_{c})-\mathrm{LAB}(\hat{\mathbf{I}}_{c})\|_{2}^{2}$, and frequency-domain loss by Fast Fourier Transforms (FFT) $\mathcal{L}_{fft}=\|\mathcal{F}(\mathbf{I}_{c})-\mathcal{F}(\hat{\mathbf{I}}_{c})\|_{2}^{2}$. Moreover, content supervision is incorporated by a CLIP loss [radford2021learning] to ensure semantic alignment, defined as:
@@ IV-B Pre-Training and Domain-specific Finetuning
The overall objective functions are as follows:
$\mathcal{L}_{0}$ $=\mathcal{L}_{r}+\mathcal{L}_{ssim}+\mathcal{L}_{color}+\mathcal{L}_{fft}+\mathcal{L}_{clip}\text{ and}$ (2)
$\mathcal{L}_{N}$ $=\mathcal{L}_{0}+\mathcal{L}_{feat}+\mathcal{L}_{percept}\text{ for }\quad(N=1,2,3,4).$ (3)
@@ IV-C Depth-aware Stylization
We design a depth-aware stylization function: DA-WCT which extends the multi-scale blending concept from the WCT (whitening and coloring transform) literature [li2017universal]. DA-WCT integrates a physics-guided waterbody estimation method for depth-guided feature blending via WCT. First, we estimate the waterbody from the style image $\mathbf{I}_{s}$ using a physics-guided method [siddique2024aquafuse]; from its depth map $\mathbf{D}_{s}$, we leverage the $5\%$ farthest pixels’ median values to estimate the waterbody background ($\mathbf{B}_{s}$) in $\mathbf{I}_{s}$. A color filter is then applied to retain only bluish-green pixels, removing red-channel influence due to differential light attenuation underwater. Let $F_{c}$ and $F_{s}$ denote the content and style feature maps extracted from the encoder, respectively. We first perform a WCT-based stylization to align the statistics of $F_{c}$ with those of $F_{s}$; the stylized feature $\tilde{F}_{cs}$ is computed as $\tilde{F}_{cs}\;=\;\mu_{s}\;+\;C_{s}^{1/2}\,C_{c}^{-1/2}\,\small(F_{c}-\mu_{c}\small)$, with $\mu_{c}$, $\mu_{s}$ being the channel-wise means and $C_{c}$, $C_{s}$ the covariance matrices of $F_{c}$ and $F_{s}$, respectively. Let $\mathbf{D}_{c}$ denote the single-channel content depth map, a global threshold $\tau\in[0,1]$ is computed using Otsu’s method [xu2011characteristic]. Next, an adaptive factor $k$ is derived by first smoothing $\mathbf{D}_{c}$ via average pooling with a $5\times 5$ kernel, computing the global mean and standard deviation as:
$\footnotesize\mu_{d}$ $=\frac{1}{H\cdot W}\sum_{x}\mathrm{avgpool}\big(\mathbf{D}_{c}(x)\big)\text{ and }$ (4)
@@
$\sigma_{d}$ $=\sqrt{\frac{1}{H\cdot W}\sum_{x}\Bigl(\mathrm{avgpool}\big(\mathbf{D}_{c}(x)\big)-\mu_{d}\Bigr)^{2}}.$ (5)
@@
With the tunable scaling factors $\tau$ and $k$, the depth weight map for each pixel is computed as:
@@
$\begin{split}w(x)&=\sigma\Bigl(-k\cdot\bigl(\mathbf{D}_{c}(x)-\tau\bigr)\Bigr)=\frac{1}{1+\exp\Bigl(k\cdot\bigl(\mathbf{D}_{c}(x)-\tau\bigr)\Bigr)}\text{.}\end{split}$ (6)
@@
In our formulation, pixels with higher depth values (i.e., farther objects) yield $w(x)\rightarrow 1$ and favor the stylized features $\tilde{F}_{cs}(x)$. In contrast, pixels with lower depth values get $w(x)\rightarrow 0$, preserving more of the original content features $F_{c}(x)$. When the depth map has low variance and is nearly flat, $\mathbf{D}_{c}\approx\tau$ which yield $w(x)\rightarrow 0.5$ across the image, so DA-WCT effectively reduces to a standard WCT-style blend. A global style strength parameter $\alpha\in[0,1]$ is applied to control the degree of the underlying stylization process as: $F_{cs}\;=\;\alpha\,F_{\text{depth}}\;+\;\bigl(1-\alpha\bigr)\,F_{c}$.
Furthermore, to capture both global and local style characteristics, our DA-WCT function is applied at multiple scales $s\in\{1.0,\,0.75,\,0.5,\,0.25\}$, with the fused features at each scale, to obtain multi-scale features $F_{\text{multi}}$. At each scale (stage) of the decoder, these fused features replace the corresponding content features with multi-scale fusion. The stylized image is then constructed as $\mathcal{D}$ as $\mathbf{I}_{cs}=\mathcal{D}\big(F_{\text{multi}}\big)$. Finally, a guided filter named GIF smoothing [he2012guided] is applied as a post-processing step to enhance the visual consistency of the final output image: $\mathbf{I}_{stylized}=\text{GIF}(\mathbf{I}_{cs})=\text{GIF}\big(\mathcal{D}(F_{\text{multi}})\big)$.
@@
Figure 5: Inference pipeline of UStyle: The content and style images and their depth maps are first processed by a ResNet-based encoder to extract multi-scale features. These are then aligned using our DA-WCT and progressively decoded via a multi-stage decoder and smoothed by a guided filter.
@@
Implementation Details. UStyle is implemented in PyTorch with OpenCV and Torch libraries. We use Adam optimizer with an initial learning rate of $10^{-4}$ for supervised learning. Training is conducted on a Linux system with $4$ GPUs and $32$ GB of RAM. At the first stage of training, the encoder-decoder network is pre-trained on the MSCOCO-2017 with $118$ K training samples for $10$ epochs. It is then fine-tuned on USOD10K underwater dataset with $10$ K examples for another $10$ epochs. The whole training and fine-tuning takes $82$ hours to complete. Once trained, the frozen graph is used for inference on a single GPU with $16$ GB of RAM; the inference pipeline is shown in Fig. 5.